OpenStudy (molly10119):

Community One is made up of four different species and has a total number of 400 individuals. Community Two is made up of 1,000 different species and has a total number of 20,000 individuals. Which community has higher species diversity a. community 1 b. community 2 c. they both have the same species diversity d. there is not enough information to determine species diversity.

alones (alones):

The fraction would be \[\frac{ 4 }{400 } \]which is equal to \[\frac{ 1 }{100 } \] and \[\frac{ 1,000 }{20,000 } \] which is equvilant to 1/20 so it would be the second one for every 1 out of every 20 species are different , so basically it should be community 2.

likeabossssssss (likeabossssssss):

welcome to os